OpenStudy (anonymous):

PLEASE HELP! the function f(t)=t^2+12t-18 represents a parabola Part A)Rewrite the function in the vertex form by completing the square. Part B) Determine the vertex and indicate whether it is a maximum or a minimum on the graph. how do you know? C)Determine the axis of symmetry for f(t)

OpenStudy (anonymous):

\[f \left( t \right)=\left( t^2+2 \times t \times 6+?^2 \right)-?^2-18\]

OpenStudy (anonymous):

I'm sorry i don't understand

OpenStudy (anonymous):

\[\left( a+b \right)^2=a^2+2*a*b+b^2\]

OpenStudy (anonymous):

(t+12)=t^2+2*t*12=12^2 .... then put it together?

OpenStudy (anonymous):

\[t^2+2t*6+6^2=\left( t+6 \right)^2\]

OpenStudy (anonymous):

t^2+12t+36=(t+6)^2 then?